OpenStudy (anonymous):

A landscaper dug a hole to plant a tree. The hole is shaped like a cylinder. If the hole is 1 foot 11 inches deep and has a diameter of 2 feet 1 inch, estimate how much dirt was removed.

OpenStudy (amistre64):

the volume of any cone shaped object is (1/3)(Base area)(height)

OpenStudy (anonymous):

idk

OpenStudy (amistre64):

we have a depth of 23 inches; and a base area of (12.5)^2pi

OpenStudy (amistre64):

(1/3)(156.25pi)(23) = your answer

OpenStudy (amistre64):

about 1197.9pi inches^3

OpenStudy (amistre64):

ack!!...helps if I read the entire question lol

OpenStudy (amistre64):

cylander..not cone

OpenStudy (anonymous):

Volume of a cylinder is Area of the Base * Height. Here, the base is a circle Area of a circle = pi r^2 So Volume = pi * r^2 * h Volume = 3.14 * 23^2 * 25 Volume 41526.5 in^3 This is the exact answer... But I noticed your problem says to estimate. So I would use Volume = 3.14 * 2 * 2 = 12.56 ft^3

OpenStudy (anonymous):

i was giving answer choices and neither are the choices

OpenStudy (amistre64):

(156.25pi)(23) = 3593.75 inches^3

OpenStudy (anonymous):

a 6.3 ft2 b.7.1 c.7.5 d.8.9

OpenStudy (amistre64):

ah...feet, i converted to inches to make life easier on me :)

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

6.3 should be it

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

i get 6.5..... 6.3 is close if you use a short form of pi

OpenStudy (anonymous):

please explain how u got that answer showing work is required (homework assignment)

OpenStudy (amistre64):

i turned the measurements into fractions: 1'11" = 1 11/12 = 23/12 this is your depth/height 2'1" = 2 1/12 = 25/12 this is your diameter, we need half of it. 25/12/2 = 25/24 is your radius base area is pi(r^2): pi (25/24)^2 = pi(625/576) base area times height is area of cylinder pi(625/576)(23/12) = pi(14375/6912) when we plug these into the calulator... we get 2.08pi = 2.08(3.14) = 6.5312 so we go with the closest answer

OpenStudy (anonymous):

thank you very much for your help